Parking in a lot costs. $2.25 for the first hour and $1.25 for each hour thereafter. A partial hour is charged the same as a full hour. What is the longest time someone can park for $9.
Answer by checkley79(3341)

2.25+1.25X=9.00
1.25X=9.00-2.25
1.25X=6.75
X=6.75/1.25
X=5.4 HOURS
1+5=6 HOURS IS THE MAX TIME FOR $9.00 OR LESS.